40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 17149     Entry time: Wed Sep 21 10:10:22 2022
Author: Yehonathan 
Type: Update 
Category: CDS 
Subject: C1SU2 crashed 

{Yehonathan, Anchal}

Came this morning to find that all the BHD optics watchdogs tripped. Watchdogs were restored but the all ADCs were stuck.

We realized that the C1SU2 model crashed.

We run /scripts/cds/restartAllModels.sh to reboot all machines. All the machines rebooted and burt restored to yesterday around 5 PM.

The optics seem to have returned to good alignment and are damped.

ELOG V3.1.3-